OpenStudy (anonymous):

URGENT Solve for y x=5y-4/1+3y

OpenStudy (anonymous):

y = x+4/5-3x

OpenStudy (anonymous):

why so??

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

@AbhimanyuPudi hw u got this

OpenStudy (anonymous):

????

OpenStudy (anonymous):

x(1+3y) = 5y-4 x+3xy = 5y-4 x+4 = 5y-3xy x+4 = y(5-3x) So, y= (x+4)/(5-3x)

OpenStudy (anonymous):

I see, you are a life safer thank you so much.